Jane wishes to invest $10,000 for 1 year.Part of it is to be invested at 5% and the rest at 6% If $560 interest is earned, how much is invested at each rate?
Answer by rfer(16322)

0.05x+0.06(10000-x)=560
0.05x+600-0.06x=560
-0.01x=-40
x=1000
10000-x=9000